# Mutation-acyclic quivers are totally proper

### v1

@@ 3. Cyclically ordered quivers
Let $Q$ be an $n$-vertex acyclic quiver. Fix a linear order $<$ of the vertices of $Q$ so that $v_{i}<v_{j}$ whenever $v_{i}\rightarrow v_{j}$. Let $U$ be the unipotent companion of the linearly ordered quiver $(Q,<)$. For a cycle $\mathcal{O}=(w_{0}-\cdots-w_{\ell}=w_{0})$ in $K_{Q}$, let $wt(\mathcal{O})=\prod_{i}|b_{w_{i}w_{i+1}}|$ if there is exactly one location $i$ with $w_{i}\leftarrow w_{i+1}$ (thus $w_{j}\rightarrow w_{j+1}$ for all $j\neq i$), and $wt(\mathcal{O})=0$ otherwise. Then the Markov invariant
@@
The $t^{n-1}$ coefficient of $\det(tU-U^{T})$ is negative of the trace of $U^{T}U^{-1}$. Let $N=I-U$, note $N$ is positive and strictly upper triangular. Thus
@@
and so ${\operatorname{tr}}(U^{T}\!U^{-1})=\sum_{k=0}^{n}{\operatorname{tr}}(U^{T}\!N^{k})$. We compute ${\operatorname{tr}}(U^{T}\!I)\!=n$, ${\operatorname{tr}}(U^{T}\!N)\!=-\sum_{v_{i}<v_{j}}\!b_{v_{i}v_{j}}^{2}$ and ${\operatorname{tr}}(U^{T}\!N^{k})=-\sum_{\mathcal{O}}wt(\mathcal{O}),$ where the sum is over the cycles $\mathcal{O}$ of length $k$. ∎
@@
Remark 3.23.
Amanda Schwartz has recently developed a combinatorial description for all of the coefficients of the Alexander polynomial $\Delta_{Q}(t)$ in the case that $Q$ is a tree quiver (that is, $K_{Q}$ is a tree) [19].
@@ 4. Technical Lemmas
Suppose that $v_{j}$ is a proper vertex in the COQ $Q$. Let $U=(u_{v_{i}v_{k}})$ be a unipotent companion of $Q$ with linear order $<$ such that $v_{i}<v_{j}$ for $v_{i}\in{\operatorname{In}}(v_{j})$. Then the matrix $M_{Q}(v_{j},-1)UM_{Q}(v_{j},-1)^{T}$ is a unipotent companion of $\mu_{v_{j}}(Q)$.
@@
We follow the proof of [11, Lemma 7.2] closely. Throughout ${\operatorname{In}}(v_{j})$ and ${\operatorname{Out}}(v_{j})$ respectively refer to the inset and outset of $v_{j}$ in $Q$.
For the mutated quiver $Q^{\prime}=\mu_{v_{j}}(Q)$ let $<^{\prime}$ be a linear order such that
@@
  • otherwise $<^{\prime}$ agrees with $<$ (so we have only changed the relative position of $v_{j}$).
Such an order exists because $v_{j}$ is proper in $Q$. Let $U^{\prime}=(u^{\prime}_{v_{i}v_{k}})$ be the unipotent companion of $Q^{\prime}$ with linear ordering $<^{\prime}$.
@@
Every case where both $v_{j}\leq^{\prime}v_{i},$ and $v_{j}\leq^{\prime}v_{k}$ appears in [11, Lemma 7.2], and the arguments therein apply. We consider the cases where $v_{j}$ appears after $v_{i}$ or $v_{k}$ in $<^{\prime}$. Note by construction of $<^{\prime}$ that if $w<^{\prime}v_{j}$ then $u_{v_{j}w}=0=n_{wv_{j}}$ and $b_{wv_{j}}=-u_{wv_{j}}.$
@@
Let $\pi$ be the permutation matrix which transforms $<$ into $<^{\prime}$. Equation (4.1) can be rewritten as $U^{\prime}=\pi^{T}M_{Q}(v_{j},-1)UM_{Q}(v_{j},-1)^{T}\pi$ (recall the matrices $E$ and $J$ defined in Definition 2.29). The claim follows. ∎
@@
Suppose that $v_{j}$ is a proper vertex in the COQ $Q$. Let $U$ be a unipotent companion of $Q$ with linear order $<$ such that $v_{j}<v_{k}$ for $v_{k}\in{\operatorname{Out}}(v_{j})$. Then the matrix $M_{Q}(v_{j},1)UM_{Q}(v_{j},1)^{T}$ is a unipotent companion of $\mu_{v_{j}}(Q)$.
@@
We reduce to Lemma 4.1 by introducing the opposite linearly ordered quiver $R^{op}$ of another linearly ordered $R$. The quiver of $R^{op}$ has the same vertices as $R$, but each arrow is reversed ($v\rightarrow w$ becomes $v\leftarrow w$) and the linear ordering is reversed ($v<w$ becomes $w<v$). It is easy to see that if $U_{R}$ is a unipotent companion of $R$, then $U_{R}^{T}$ is a unipotent companion of $R^{op}$ (note that $U_{R}^{T}$ is upper triangular in the reversed linear order). We also have $M_{R}(v_{j},1)=M_{R^{{op}}}(v_{j},-1)$.
@@
This matrix is upper triangular with the linear order $v_{4}<v_{2}<v_{1}<v_{3}$ (thus we have swapped the $2$nd and $3$rd rows and columns), in agreement with Lemma 4.1.
@@
This theorem generalizes an example provided in [21, Figure 1].
@@
Suppose $Q$ is mutation-acyclic, and $v$ is a vertex in $Q$. Partition the other vertices of $Q$ into three sets ${\operatorname{Out}}(v),{\operatorname{In}}(v)$ and the set of vertices $S$ which are not adjacent to $v$. Consider the subquivers supported by ${\operatorname{Out}}(v)\cup S$ (resp. ${\operatorname{In}}(v)\cup S$). Essentially the same argument as in Theorem 4.4 implies that every oriented cycle in these subquivers contains at most one vertex (and thus no arrows) in ${\operatorname{Out}}(v)$ (resp. ${\operatorname{In}}(v)$). Thus, for example, the quiver in Figure 9 is not mutation-acyclic.
@@
Figure 9. A quiver on $6$ vertices which does not have an admissible quasi-Cartan companion, and so cannot be mutation-acyclic. The subquiver supported by $v_{1},\ldots,v_{5}$ is an oriented cycle, contains two vertices in ${\operatorname{Out}}(v_{6})$, but no vertices in ${\operatorname{In}}(v_{6})$.
@@
Fix a quiver $Q$ and two linear orderings of its vertices <,<,< , and $<^{\prime}$. Let $U=(u_{vw})$ and $U^{\prime}=(u^{\prime}_{vw})$ be the unipotent companions of the linearly ordered quivers $(Q,<)$ and $(Q,<^{\prime})$ respectively. If $u_{vw}=u^{\prime}_{vw}$ for every pair of vertices $v,w$, then the COQs associated to $(Q,<)$ and $(Q,<^{\prime})$ are wiggle equivalent.
@@
By Theorem 3.11, it suffices to check that both COQs have the same winding numbers. From (3.1), it suffices to identify the arrows which “wrap around” the respective linear orderings (and, if they differ, their orientations). For any two vertices $v,w$ we have $v<w$ if and only if $u_{vw}\neq 0$. Thus the same arrows contribute regardless of if we use the linear ordering $<$ or $<^{\prime}$, and so the winding numbers agree. ∎
@@
Let $Q_{0}$ be the acyclic quiver of type $E_{6}$ shown (as a COQ) in Figure 10, and let $t\in\mathbb{T}_{n}$ be the vertex such that
@@
Two quivers on six vertices v1, v2, v3, v4, v5, v6. On the left, a tree quiver with an arrow from v1 to v2, v2 to v3, v3 to v4, v4 to v5, and v3 to v6. On the right, we have an arrow from v1 to v4, v4 to v2, v2 to v6, v6 to v1, v6 to v3, v3 to v2, v2 to v5, and v5 to v3.
@@
The directed graphs associated to vertices v1, v2, and v3. In each, vertices v3 and v6 are colored green, while v1, v2, v4, and v5 are colored red. Each has six vertices v1, v2, v3, v4, v5, v6, and all have edges from v1 to v4, v4 to v2, v2 to v5, v6 to v3, v6 to v2, v6 to v1, v3 to v2, and v3 to v5. The graph associated to v1 has no other edges. The graph associated to v2 has an additional edge from v6 to v3 which is labeled v2. The graph associated to v3 has an additional edge from v2 to v5 labeled v3.
The graphs associated to vertices v4, v5, and v6. The graphs have the same vertices and arrows as v1 in the previous figure, with the only additional edge appearing in the graph associated to v6. The new edge is from v1 to v2, and is labeled v6.
@@
A linear extension of $L_{v}(Q_{t})$ is a linear order $<$ on the vertices such that $v_{i}<v_{j}$ whenever $v_{i}\rightarrow v_{j}$ (it does not matter whether the arrow is labeled $v$). Let $\sigma_{v}$ be a cyclic ordering associated to a linear extension of $L_{v}(Q_{t})$.
@@
The matrix $U_{t}=(u_{v_{i}v_{j};t})$ is upper triangular when written with a linear extension $<$ of $L_{v}(Q_{t})$. Thus $U_{t}$ is a unipotent companion of $(Q_{t},\sigma_{v})$.
@@
It suffices to check that each vertex $q$ is proper in the COQ $(Q_{t},\sigma_{q})$ by Lemma 5.10. Let $<$ be a linear extension of $L_{q}(Q_{t})$. We consider all oriented paths ${p\rightarrow q\rightarrow r}$ in cases, based on the color (green or red) of $p,q,$ and $r$ in $Q_{t}$. As noted in Lemma 5.5, the subquiver of all red or all green vertices is acyclic. We have chosen our order $<$ to be compatible with these acyclic subquivers. Thus if all three of $p,q,r$ are green (resp., red) then $p\rightarrow q\rightarrow r$ is a right turn at $q$. A similar argument applies if $p$ or $r$ is a different color than the other two vertices.
This leaves us with the alternating patterns: red-green-red or green-red-green. By construction of $<$, if $p,r$ are both a different color than $q$, then there is an arrow $r\stackrel{{\scriptstyle q}}{{\rightarrow}}p$ in $L_{q}(Q_{t})$, and so $r<p$. Thus the path $p\rightarrow q\rightarrow r$ is again a right turn at $q$. ∎
@@
Continuing with Examples 5.2 and 5.7, Figure 12 shows the COQ $(Q_{t},\sigma_{v_{1}})$. Note that every oriented path of length $2$ is a right turn.
@@
Fix an arbitrary $t^{\prime}\stackrel{{\scriptstyle i}}{{\text{---}}}t$. It suffices to show that the COQ $(\mu_{v_{i}}(Q_{t}),\sigma_{t^{\prime}})$ is in the wiggle equivalence class $\mu_{v_{i}}((Q_{t},\sigma_{t}))$. Let $<$ be a linear extension of $L_{v_{i}}(Q_{t})$. Say that $v_{i}$ is green (resp., red) in $Q_{t}$. For every vertex $w\in{\operatorname{Out}}(v_{i})$ (resp., ${\operatorname{In}}(v_{i})$), we have $v_{i}<w$ (resp. $w<v_{i}$). By Proposition 2.30, we have
@@
for $\varepsilon=1$ (resp., $-1$). Thus
@@ 6. Applications
Theorem 6.1 ([11, Theorem 11.3, Remark 11.11]).
@@
As an illustration, we compute some mutation invariants for a few families. Let $\tilde{A}(r,\ell)$ denote a quiver with $n=r+\ell$ vertices $v_{1},\ldots,v_{n}$ which has $r$ locations with a single arrow $v_{i}\rightarrow v_{i+1}$ and $\ell$ locations with a single arrow $v_{i}\leftarrow v_{i+1}$ (see [13, Example 6.9]). While there are many ways to place the arrows, all are related by sink and/or source mutations. Thus we may assume that $\tilde{A}(r,\ell)$ has arrows
$v_{1}\rightarrow v_{2}\rightarrow\cdots\rightarrow v_{r}\leftarrow v_{r+1}\leftarrow\cdots\leftarrow v_{n}\leftarrow v_{1}.$
@@
$v_{1}\stackrel{{\scriptstyle a_{1}}}{{\rightarrow}}v_{2}\stackrel{{\scriptstyle a_{2}}}{{\rightarrow}}\cdots\stackrel{{\scriptstyle a_{r-1}}}{{\rightarrow}}v_{r}\stackrel{{\scriptstyle a_{r}}}{{\leftarrow}}v_{r+1}\stackrel{{\scriptstyle a_{r+1}}}{{\leftarrow}}\cdots\stackrel{{\scriptstyle a_{n-1}}}{{\leftarrow}}v_{n}\stackrel{{\scriptstyle a_{n}}}{{\leftarrow}}v_{1}.$
@@
Let $Q$ be a connected, totally proper and mutation-acyclic COQ on $n$ vertices. Then the Markov invariant $M_{Q}\geq n-1$. Further, if $M_{Q}=n-1$ then $Q$ is mutation equivalent to a quiver whose underlying undirected graph is a tree (with no parallel arrows).
@@
It suffices to consider acyclic COQs $Q$ with vertices $v_{1},\ldots,v_{n}$. By Proposition 3.21, $\sum_{i<j}b_{v_{i}v_{j}}^{2}\leq M_{Q}.$ Recall the Frobenius norm $||B_{Q}||_{F}=\big{(}\sum_{i,j}b_{v_{i}v_{j}}^{2}\big{)}^{1/2}$ of $B_{Q}$ (see also the Hilbert-Schmidt norm or Schur norm), thus $||B_{Q}||_{F}^{2}\leq 2M_{Q}.$
As $\det(B_{Q})=0$ if $n$ is odd, we assume $n$ is even. As $B_{Q}$ is skew-symmetric, its eigenvalues are all imaginary and come in signed pairs, say they have norms $|\lambda_{1}|,\ldots,|\lambda_{n/2}|$. Let $\lambda=\max(|\lambda_{j}|)$. Then $\det(B_{Q})\leq\lambda^{n}$. As $B_{Q}$ is skew-symmetric the spectral norm $\rho(B_{Q})=\lambda$. It is a classical fact that the spectral norm is a lower bound on the Frobenius norm: $\lambda\leq||B_{Q}||_{F}$. Therefore we have
$\det(B_{Q})\leq\lambda^{n}\leq||B_{Q}||_{F}^{n}\leq(2M_{Q})^{n/2}.\qed$
Corollaries 6.6 and 6.7 give short proofs that certain quivers are not mutation-acyclic.
@@
A COQ with cyclic ordering (v1, v2, v3, v4) and a1 arrows from v1 to v2, a2 arrows from v2 to v3, a3 arrows from v3 to v4, and a4 arrows from v4 to v1.
Some such quivers are mutation-acyclic, for example when $a_{i}=1$ for all $i$ then $C(4,0)$ has type $D_{4}$. Others are not. Warkentin showed in [26][Example 6.11] that if $a_{1}=a_{3}\geq 2$ and $a_{2}=a_{4}\geq 2$ then this quiver is not mutation-acyclic (by a precise description of the mutation class). The same can be observed by Corollary 6.6, noting that $M_{C(4,0)}\leq 0$ for such quivers.
@@
Consider the quivers associated with the elliptic root system of double extended type $E_{7}$ or $E_{8}$, denoted $E_{7}^{(1,1)}$ or $E_{8}^{(1,1)}$ and shown in Figure 14. (Note that the quiver in Figure 1 is mutation equivalent to $E_{8}^{(1,1)}$.) These quivers have finite mutation classes, and can be shown to be totally proper by exhaustive computation. (We emphasize the exhaustion; $E_{8}^{(1,1)}$ has $5739$ quivers, up to isomorphism, in its class.) These quivers are not mutation-acyclic. This has been shown by brute force, or by representation theory [17][Section 2.3]. We sketch a new proof.
@@
So it suffices to check if $\Delta_{E_{7}^{(1,1)}}(t)$ and $\Delta_{E_{8}^{(1,1)}}(t)$ appear in these lists. A (still tedious, but straightforward and human verifiable) computation shows that they are not. The computation can be further simplified using the work of Amanda Schwartz, see Remark 3.23.

### v2

@@ 3. Cyclically ordered quivers
Let $Q$ be an $n$-vertex acyclic quiver. Fix a linear order << of the vertices of $Q$ so that $v_{i}<v_{j}$ whenever $v_{i}\rightarrow v_{j}$. Let $U$ be the unipotent companion of the linearly ordered quiver $(Q,<)$. For a cycle $\mathcal{O}=(w_{0}-\cdots-w_{\ell}=w_{0})$ in $K_{Q}$, let $wt(\mathcal{O})=\prod_{i}|b_{w_{i}w_{i+1}}|$ if there is exactly one location $i$ with $w_{i}\leftarrow w_{i+1}$ (thus $w_{j}\rightarrow w_{j+1}$ for all $j\neq i$), and $wt(\mathcal{O})=0$ otherwise. Then the Markov invariant
@@
The $t^{n-1}$ coefficient of $\det(tU-U^{T})$ is negative of the trace of $U^{T}U^{-1}$. Let N=IUN=I-U, note $N$ is positive and strictly upper triangular. Thus
@@
and so ${\operatorname{tr}}(U^{T}\!U^{-1})=\sum_{k=0}^{n}{\operatorname{tr}}(U^{T}\!N^{k})$. We compute ${\operatorname{tr}}(U^{T}\!I)\!=n$, ${\operatorname{tr}}(U^{T}\!N)\!=-\sum_{v_{i}<v_{j}}\!b_{v_{i}v_{j}}^{2}$ and ${\operatorname{tr}}(U^{T}\!N^{k})=-\sum_{\mathcal{O}}wt(\mathcal{O}),$ where the sum is over the cycles $\mathcal{O}$ of length $k$. (Recall that cycles are considered up to cyclic shifts.) ∎
@@ 4. Technical Lemmas
Suppose that $v_{j}$ is a proper vertex in the COQ $Q$. Let $U=(u_{v_{i}v_{k}})$ be a unipotent companion of $Q$ with linear order << such that $v_{i}<v_{j}$ for $v_{i}\in{\operatorname{In}}(v_{j})$. Then the matrix $M_{Q}(v_{j},-1)UM_{Q}(v_{j},-1)^{T}$ is a unipotent companion of $\mu_{v_{j}}(Q)$.
@@
We follow the proof of [13, Lemma 7.2] closely. Throughout ${\operatorname{In}}(v_{j})$ and ${\operatorname{Out}}(v_{j})$ respectively refer to the inset and outset of $v_{j}$ in $Q$.
For the mutated quiver $Q^{\prime}=\mu_{v_{j}}(Q)$ let $^{\prime}$ be a linear order such that
@@
  • otherwise $^{\prime}$ agrees with << (so we have only changed the relative position of $v_{j}$).
Such an order exists because $v_{j}$ is proper in $Q$. Let $U^{\prime}=(u^{\prime}_{v_{i}v_{k}})$ be the unipotent companion of $Q^{\prime}$ with linear ordering $^{\prime}$.
@@
Every case where both $v_{j}\leq^{\prime}v_{i},$ and $v_{j}\leq^{\prime}v_{k}$ appears in [13, Lemma 7.2], and the arguments therein apply. We consider the cases where $v_{j}$ appears after $v_{i}$ or $v_{k}$ in $^{\prime}$. Note by construction of $^{\prime}$ that if $w<^{\prime}v_{j}$ then $u_{v_{j}w}=0=n_{wv_{j}}$ and $b_{wv_{j}}=-u_{wv_{j}}.$
@@
Let $\pi$ be the permutation matrix which transforms << into $^{\prime}$. Equation (4.1) can be rewritten as $U^{\prime}=\pi^{T}M_{Q}(v_{j},-1)UM_{Q}(v_{j},-1)^{T}\pi$ (recall the matrices $E$ and $J$ defined in Definition 2.29). The claim follows. ∎
@@
Suppose that $v_{j}$ is a proper vertex in the COQ $Q$. Let $U$ be a unipotent companion of $Q$ with linear order << such that $v_{j}<v_{k}$ for $v_{k}\in{\operatorname{Out}}(v_{j})$. Then the matrix $M_{Q}(v_{j},1)UM_{Q}(v_{j},1)^{T}$ is a unipotent companion of $\mu_{v_{j}}(Q)$.
@@
We reduce to Lemma 4.1 by introducing the opposite linearly ordered quiver $R^{op}$ of another linearly ordered $R$. The quiver of $R^{op}$ has the same vertices as $R$, but each arrow is reversed ($v\rightarrow w$ becomes $v\leftarrow w$) and the linear ordering is reversed ($v<w$ becomes $v>w$). It is easy to see that if $U_{R}$ is a unipotent companion of $R$, then $U_{R}^{T}$ is a unipotent companion of $R^{op}$ (note that $U_{R}^{T}$ is upper triangular in the reversed linear order). We also have $M_{R}(v_{j},1)=M_{R^{{op}}}(v_{j},-1)$.
@@
This matrix is upper triangular with the linear order $v_{4}<v_{2}<v_{1}<v_{3}$ (thus we have swapped the 22nd and 33rd rows and columns), in agreement with Lemma 4.1.
@@
This theorem generalizes an example provided in [24, Figure 1].
@@
Suppose $Q$ is mutation-acyclic, and $v$ is a vertex in $Q$. Partition the other vertices of $Q$ into three sets ${\operatorname{Out}}(v),{\operatorname{In}}(v)$ and the set of vertices $S$ which are not adjacent to $v$. Consider the subquiver supported by ${\operatorname{Out}}(v)\cup S$ (resp. ${\operatorname{In}}(v)\cup S$). Essentially the same argument as in Theorem 4.4 implies that every pair of vertices in ${\operatorname{Out}}(v)$ (resp. ${\operatorname{In}}(v)$) on a chordless oriented cycle in ${\operatorname{Out}}(v)\cup S$ (resp. ${\operatorname{In}}(v)\cup S$) must be adjacent. Thus, for example, the quiver in Figure 9 is not mutation-acyclic.
@@
Figure 9. A quiver on $6$ vertices which has no admissible quasi-Cartan companion, and so is not mutation-acyclic. The subquiver supported by $v_{1},\ldots,v_{5}$ is an oriented cycle, contains two non-adjacent vertices in ${\operatorname{Out}}(v_{6})$, but no vertices in ${\operatorname{In}}(v_{6})$.
@@
Fix a quiver $Q$ and two linear orderings of its vertices <,<, and $^{\prime}$. Let $U=(u_{vw})$ and $U^{\prime}=(u^{\prime}_{vw})$ be the unipotent companions of the linearly ordered quivers $(Q,<)$ and $(Q,<^{\prime})$ respectively. If $u_{vw}=u^{\prime}_{vw}$ for every pair of vertices $v,w$, then the COQs associated to $(Q,<)$ and $(Q,<^{\prime})$ are wiggle equivalent.
@@
By Theorem 3.11, it suffices to check that both COQs have the same winding numbers. From (3.1), it suffices to identify the arrows which “wrap around” the respective linear orderings (and, if they differ, their orientations). For any two vertices $v,w$ we have $v<w$ if and only if $u_{vw}\neq 0$. Thus the same arrows contribute regardless of if we use the linear ordering << or $^{\prime}$, and so the winding numbers agree. ∎
@@
Let $Q_{0}$ be the acyclic quiver of type $E_{6}$ shown in Figure 10, and let $t\in\mathbb{T}_{n}$ be the vertex such that
@@
A tree quiver on six vertices v1, v2, v3, v4, v5, v6, with an arrow from v1 to v2, v2 to v3, v3 to v4, v4 to v5, and v3 to v6.
A tree quiver on six vertices v1, v2, v3, v4, v5, v6, with an arrow from v1 to v4, v4 to v2, v2 to v6, v6 to v1, v6 to v3, v3 to v2, v2 to v5, and v5 to v3.
@@
The directed graphs associated to vertices v1, v2, and v3. In each, vertices v3 and v6 are colored green, while v1, v2, v4, and v5 are colored red. Each has six vertices v1, v2, v3, v4, v5, v6, and all have edges from v1 to v4, v4 to v2, v2 to v5, v6 to v3, v6 to v2, v6 to v1, v3 to v2, and v3 to v5. The graph associated to v1 has no other edges.
The graph associated to v2 has an additional edge from v6 to v3 which is labeled v2.
 The graph associated to v3 has an additional edge from v2 to v5 labeled v3.
The graph associated to v4 is identical to that of v1.
The graph associated to v5 is identical to that of v1.
The graph associated to v6 is identical to that of v1.
@@
A linear extension of $L_{v}(Q_{t})$ is a linear order << on the vertices such that $v_{i}<v_{j}$ whenever $v_{i}\rightarrow v_{j}$ (it does not matter whether the arrow is labeled $v$). Let $\sigma_{v}$ be a cyclic ordering associated to a linear extension of $L_{v}(Q_{t})$.
@@
The matrix $U_{t}=(u_{v_{i}v_{j};t})$ is upper triangular when written with a linear extension << of $L_{v}(Q_{t})$. Thus $U_{t}$ is a unipotent companion of $(Q_{t},\sigma_{v})$.
@@
It suffices to check that each vertex $q$ is proper in the COQ $(Q_{t},\sigma_{q})$ by Lemma 5.10. Let << be a linear extension of $L_{q}(Q_{t})$. We consider all oriented paths $p\rightarrow q\rightarrow r$} in cases, based on the color (green or red) of $p,q,$ and $r$ in $Q_{t}$. As noted in Lemma 5.5, the subquiver of all red or all green vertices is acyclic. We have chosen our order << to be compatible with these acyclic subquivers. Thus if all three of $p,q,r$ are green (resp., red) then $p\rightarrow q\rightarrow r$ is a right turn at $q$. A similar argument applies if $p$ or $r$ is a different color than the other two vertices.
This leaves us with the alternating patterns: red-green-red or green-red-green. By construction of <<, if $p,r$ are both a different color than $q$, then there is an arrow $r\stackrel{{\scriptstyle q}}{{\rightarrow}}p$ in $L_{q}(Q_{t})$, and so $r<p$. Thus the path $p\rightarrow q\rightarrow r$ is again a right turn at $q$. ∎
@@
Continuing with Examples 5.2 and 5.7, Figure 12 shows the COQ $(Q_{t},\sigma_{v_{1}})$. Note that every oriented path of length 22 is a right turn.
@@
Fix an arbitrary $t^{\prime}\stackrel{{\scriptstyle i}}{{\text{---}}}t$. It suffices to show that the COQ $(\mu_{v_{i}}(Q_{t}),\sigma_{t^{\prime}})$ is in the wiggle equivalence class $\mu_{v_{i}}((Q_{t},\sigma_{t}))$. Let << be a linear extension of $L_{v_{i}}(Q_{t})$. Say that $v_{i}$ is green (resp., red) in $Q_{t}$. For every vertex $w\in{\operatorname{Out}}(v_{i})$ (resp., ${\operatorname{In}}(v_{i})$), we have $v_{i}<w$ (resp. $w<v_{i}$). By Proposition 2.30, we have
@@
for $\varepsilon=1$ (resp., 1-1). Thus
@@ 6. Applications
Theorem 6.1 ([13, Theorem 14.3, Algorithm 14.12]).
@@
As an illustration, we compute some mutation invariants for a few families. Let $\tilde{A}(r,\ell)$ denote a quiver with $n=r+\ell$ vertices $v_{1},\ldots,v_{n}$ which has $r$ locations with a single arrow $v_{i}\rightarrow v_{i+1}$ and $\ell$ locations with a single arrow $v_{i}\leftarrow v_{i+1}$ (see [15, Example 6.9]). While there are many ways to place the arrows, all are related by sink and/or source mutations. Thus we may assume that $\tilde{A}(r,\ell)$ has arrows
$v_{1}\rightarrow v_{2}\rightarrow\cdots\rightarrow v_{r}\rightarrow v_{r+1}\leftarrow\cdots\leftarrow v_{n}\leftarrow v_{1}.$
@@
$v_{1}\stackrel{{\scriptstyle a_{1}}}{{\rightarrow}}v_{2}\stackrel{{\scriptstyle a_{2}}}{{\rightarrow}}\cdots\stackrel{{\scriptstyle a_{r-1}}}{{\rightarrow}}v_{r}\stackrel{{\scriptstyle a_{r}}}{{\rightarrow}}v_{r+1}\stackrel{{\scriptstyle a_{r+1}}}{{\leftarrow}}\cdots\stackrel{{\scriptstyle a_{n-1}}}{{\leftarrow}}v_{n}\stackrel{{\scriptstyle a_{n}}}{{\leftarrow}}v_{1}.$
@@
If $Q$ is a totally proper mutation-acyclic COQ on $n$ vertices then $M_{Q}\geq 0$. If, additionally, $Q$ is connected then the Markov invariant $M_{Q}\geq n-1$.
Corollary 6.7.
For a fixed integer $k$, there are only finitely many acyclic quivers $Q$ with $M_{Q}=k$ (and hence only finitely many acyclic quivers with the same Alexander polynomial). In particular, if $M_{Q}=n-1$ and $Q$ is connected then $Q$ is mutation equivalent to a quiver whose underlying undirected graph is a tree (with no parallel arrows).
Note that $M_{Q}$ is monotone in all of the entries of $B_{Q}$ for an acyclic quiver $Q$. Therefore we can enumerate the acyclic quivers with $M_{Q}=k$ by searching all acyclic quivers with weights at most, say, $\sqrt{k}$ for matching Markov invariants.
@@
It suffices to consider acyclic COQs $Q$ with vertices $v_{1},\ldots,v_{n}$. By Proposition 3.21, $\sum_{i<j}b_{v_{i}v_{j}}^{2}\leq M_{Q}.$ Recall the Frobenius norm $B_{Q}||_{F}=\big(\sum_{i,j}b_{v_{i}v_{j}}^{2}\big)^{1/2}$ of $B_{Q}$ (see also the Hilbert-Schmidt norm or Schur norm), thus $B_{Q}||_{F}^{2}\leq 2M_{Q}.$
[…]
Corollaries 6.66.7 and 6.8 give short proofs that certain quivers are not mutation-acyclic.
@@
A COQ with cyclic ordering (v1,v2,v3,v4) and a(i) arrows from v(i) to v(i+1) (indicies mod 4).
Some such quivers are mutation-acyclic, for example when $a_{i}=1$ for all $i$ then $C(4,0)$ has type $D_{4}$. Others are not. Warkentin showed in [29, Example 6.11] that if $a_{1}=a_{3}\geq 2$ and $a_{2}=a_{4}\geq 2$ then this quiver is not mutation-acyclic (by a precise description of the mutation class). The same can be observed by Corollary 6.6, noting that $M_{C(4,0)}\leq 0$ for such quivers.
@@
As an application of Corollary 6.7, we enumerate all of the acyclic $4$-vertex quivers with $M_{Q}\leq 6$ (up to isomorphism and mutation equivalence). Because $M_{Q}\geq\sum_{i<j}b_{ij}^{2}$, we may assume $b_{ij}\leq 2.$ If $M_{Q}<3$ then $Q$ must be disconnected. There is only one quiver with 0 or $1$ arrows. There are two quivers with $M_{Q}=2$, corresponding to type $A_{3}$ (with an isolated vertex), and the union of two copies of $A_{2}$. Up to mutation equivalence, there are two tree quivers on $4$-vertecies, corresponding to types $A_{4},D_{4}$, both of which have $M_{Q}=3$. The only mutation class of connected acyclic quivers with $M_{Q}=4$ is $\tilde{A}(2,2)$. There are also the disconnected quivers corresponding to $v_{1}\stackrel{{\scriptstyle 2}}{{\rightarrow}}v_{2}$ and $\tilde{A}(2,1)$ (with all other vertices isolated). There are two mutation-classes of connected acyclic quivers $Q$ with $M_{Q}=5$, as well as two classes of disconnected quivers. Representitives are shown below:
Four COQs, each with cyclic ordering (v1, v2, v3, v4). The leftmost has arrows from v1 to v2, v2 to v3, v3 to v4, and v1 to v4. The next has an arrow from v1 to v2, v2 to v3, v3 to v4, and v1 to v3. The next has 2 arrows from v1 to v2, and one arrow from v3 to v4. The rightmost has 2 arrows from v1 to v2, and one arrow from v3 to v4.
Any 4-vertex acyclic quiver with $M_{Q}=6$ must be connected. A quiver from each mutation-acyclic mutation-class with $M_{Q}=6$ are shown below.
Three COQs, each with cyclic ordering (v1,v2,v3,v4). The leftmost has 2 arrows from v1 to v2, and just one arrow from v2 to v3, and v3 to v4. The middle COQ has one arrow from v1 to v2 and v3 to v4, and two arrows from v2 to v3. The last has 2 arrows from v1 to v2, and just one arrow from v1 to v3 and v1 to v4.
Thus the complete list of possible Alexander polynomials of mutation-acyclic quivers with $M_{Q}\leq 6$} is:
[…]
There is only one polynomial with a $4t(t-1)^{2}$ summand because all three mutation-classes with $M_{Q}=4$ have the same Alexander polynomial. (As do two of the mutation-classes with $M_{Q}=5$.)
Example 6.13.
Consider the quivers associated with the elliptic root system of double extended type $E_{7}$ or $E_{8}$, denoted $E_{7}^{(1,1)}$ or $E_{8}^{(1,1)}$ and shown in Figure 14. (Note that the quiver in Figure 1 is mutation equivalent to $E_{8}^{(1,1)}$.) These quivers have finite mutation classes, and can be shown to be totally proper by exhaustive computation. (We emphasize the exhaustion; $E_{8}^{(1,1)}$ has $5739$ quivers, up to isomorphism, in its class.) These quivers are not mutation-acyclic. This has been shown by brute force, or by representation theory [19, Section 2.3]. We sketch a new proof.
@@
So it suffices to check if
$\Delta_{E_{7}^{(1,1)}}(t)$ $=t^{9}-t^{8}-2t^{5}+2t^{4}+t-1,\text{ and}$
$\Delta_{E_{8}^{(1,1)}}(t)$ $=t^{10}-t^{9}+t^{7}-t^{6}-t^{4}+t^{3}-t+1$
appear in these lists. A (still tedious, but straightforward and human verifiable) computation shows that they are not. The computation can be further simplified using the work of Amanda Schwartz [21].